OpenStudy (anonymous):

For an arts and crafts project, a student cut a triangle out of a rectangular piece of paper. If the triangle has a base of 6 inches and a height of 8 inches, and the rectangle has a length of 11 inches and a width of 12 inches, what is the area of the leftover paper after the triangle is cut out?

ILovePuppiesLol (ilovepuppieslol):

Welcome to open study!

OpenStudy (anonymous):

For an arts and crafts project, a student cut a triangle out of a rectangular piece of paper. If the triangle has a base of 6 inches and a height of 8 inches, and the rectangle has a length of 11 inches and a width of 12 inches, what is the area of the leftover paper after the triangle is cut out?

Directrix (directrix):

Calculate the area of the triangle and rectangle. Subtract the area of the triangle from the area of the rectangle.

Directrix (directrix):

Area of a rectangle is L*w so' Where l is length and w is width. A = l*w A = 11*12 A = ?

Directrix (directrix):

Area of a triangle is 1/2b * h A = 1/2b * h where b is base and h is height A = 1/2 * 6 * 8 A = ?

Directrix (directrix):

Area of leftover paper is: Area of Rectangle - Area of Triangle @Kceeingram123

Directrix (directrix):

Answer Options: ANSWERS: a. 108 in2 b. 150 in2 c. 174 in2 d. 24 in2